An impulse of 10 kilogram-meter per second acting on an object whose mass is 5 kilogram will cause a change in the objects velocity of: find the solution
1
Expert's answer
2019-06-03T10:39:38-0400

Impulse acting on an object is the change in momentum of an object


"\Delta p=\Delta \left( mv \right)"

Since the mass m does not change, then impulse of object is

"\Delta p=m\Delta v"

Dividing both sides of the equation by m, we get the change in velocity


"\Delta v=\frac{\Delta p}{m}"

Substituting "\Delta p=10\text{ }kg\text{-}meter\text{ }per\text{ }second" and "m=5\text{ }kg" we get change in the objects velocity


"\Delta v=\frac{10\,\,kg\text{-}meter\text{ }per\text{ }second}{5\,kg}=2\,meter\text{ }per\text{ }second"